OpenStudy (anonymous):

The sum of three consecutive numbers is 216. What are the three numbers and how do you come up with them?

OpenStudy (anonymous):

First consider the mathematical view of the problem, set it up with algebra. x+(x+1)+(x+2) = 216 Sum the terms: 3x+3 = 216 Solve for x: 3x = 213 x = 71 And the two other numbers: x+1 = 72 x+2 = 73

OpenStudy (anonymous):

Thank you.. You up for another?

OpenStudy (anonymous):

Sure, but try it yourself first :) Always try, see what you get, post your results and it's easier to help and you learn more.

OpenStudy (anonymous):

Jason had to submit five animation projects as per his syllabus in the final semesters. He scored 88, 79, 76, and 90 points out of 100 in four projects. What should be his score in the fifth project so that the average of all the projects is at least 83?